OpenStudy (anonymous):

lim as y-->3 of (y^2+y-12)/(y^3-10y+3). I just can't figure out how to factor this thing, really.

OpenStudy (helder_edwin):

u got \[ \large \lim_{y\to3}\frac{y^2+y-12}{y^3-10y+3}=\frac{0}{0} \] right?

OpenStudy (helder_edwin):

this means that the factor x-3 is present in both the numerator and the denominator

OpenStudy (helder_edwin):

u can then use synthetic division to factor them both

OpenStudy (helder_edwin):

@CaptainMathNoob ?? r u there?

OpenStudy (anonymous):

Does it explicitly say to factor, or are you familiar w/ L'Hospitals? Can you do a synthetic division?

OpenStudy (anonymous):

Yeah I'm here, sorry. This is a Calc 1 class and we haven't gotten to L'Hospitals yet.

OpenStudy (anonymous):

Ok, so do you know how to do a synthetic division?

OpenStudy (anonymous):

I'm vaguely familiar with it. Guess I should go watch a youtube video, eh?

OpenStudy (anonymous):

I'll work the numerator, but yes. You should go revise that.

OpenStudy (anonymous):